Supposed a certain game is fair and costs $10 if you lose and has a net payoff of $9 if you win. The only possible outcomes of the game are winning and losing.

What is the probability of winning?
Answer #1

the expected value of a game to be 0.

let the probability of winning is x,then the probability of losing is (1-x);

for the expected value to be 0, we need to have

9x +(-10)(1-x)=0

9x+10x=9

X=9/19

Probability =9/19=0.474